OpenStudy (anonymous):

this one is tricky. help please! y''-4y'+3y=2cosx+4sinx i'm having trouble getting the Yp!

OpenStudy (anonymous):

You can try substituting y = asinx + bcosx

OpenStudy (anonymous):

why would that help. I'm in diff eq as well and I wasn't thinking about that.

OpenStudy (anonymous):

its called th method of undetermined coefficients

OpenStudy (anonymous):

put it on the left side....compare with coeffs on the right fr etermining a a and b

OpenStudy (anonymous):

isn't it y=a cos x + b sinx + cx cos x + dx sin x

OpenStudy (anonymous):

that is for a special case i think

OpenStudy (anonymous):

what are the roots anyway? +i, -i aren't they like, twice? then it will be case 4?

OpenStudy (anonymous):

r^2 - 4r +3 = 0 (r - 3)(r-1) r= 3, 1

OpenStudy (anonymous):

that is for the homo part

OpenStudy (anonymous):

im telling you take y=asinx + bcosx substitute y'' and y' and y in the above eqn and compare LHS and RHS ull get 2 eqns in a and b determine a and b and you have your soln

OpenStudy (anonymous):

i did that, i always get 0 for B. it's not supposed to be the answer.

OpenStudy (anonymous):

im getting the answer as b=1, a=0 which means y=cos x y=cos x satisfies the above eqns

OpenStudy (anonymous):

oh sorry. i looked at the wrong answer on my book. i am sorry. Our answer is right. Thank you!

OpenStudy (anonymous):

Your welcome